# Learning From Failure: Integrating Negative Examples when Fine-tuning Large Language Models as Agents

### v1

@@ Abstract
Large language models (LLMs) have achieved success in acting as agents, which interact with environments through tools like search engines. However, LLMs are not optimized specifically for tool use during training or alignment, limiting their effectiveness as agents. To resolve this problem, previous work has collected interaction trajectories between GPT-4 and environments, and fine-tuned smaller models with them. As part of this, the standard approach has been to simply discard trajectories that do not finish the task successfully, which, on the one hand, leads to a significant waste of data and resources, and on the other hand, has the potential to limit the possible optimization paths during fine-tuning. In this paper, we contend that large language models can learn from failures through appropriate data cleaning and fine-tuning strategies. We conduct experiments on mathematical reasoning, multi-hop question answering, and strategic question answering tasks. Experimental results demonstrate that compared to solely using positive examples, incorporating negative examples enhances model performance by a large margin. [^1] Code and data are available at: https://github.com/Reason-Wang/NAT.
@@ 1 Introduction
An agent is a model which has the ability to interact with an environment, make decisions, and achieve predefined goals (Wooldridge, 1999). Early work used rule-based or template-based systems to complete tasks in narrow and specialized domains (Green Jr et al., 1961; Weizenbaum, 1966). Recent work has taken powerful LLMs such as GPT-4 (Achiam et al., 2023) ans used them as the core of an agent system to process information and make decisions (Gravitas, 2024; Yoheinakajima, 2024). This line of work has resulted in agent systems that are able to do much more complex and general tasks.
@@
However, these models rely on closed models through paid APIs, raising concerns about cost, latency, and reproducibility. Additionally, existing LLMs were not developed for agent use cases (e.g., generating actions), and few-shot prompting offers only limited learning support (Chen et al., 2023).
@@
Subsequent work has explored fine-tuning LLMs as agents, typically in three stages: data collection, fine-tuning, and inference (Chen et al., 2023; Zeng et al., 2023; Yin et al., 2023; Qiao et al., 2024). At the data collection stage, a powerful LLM such as GPT-4 is employed to interact with the environment, and the LLM-generated outputs and environment observations are collected as trajectories. In the fine-tuning stage, smaller models are fine-tuned using only successful trajectories. The fine-tuned models then serve as the agent’s core during inference, demonstrating enhanced tool-using and decision-making capabilities, sometimes even surpassing the performance of the original LLM.
@@
To ensure the agent is being optimized appropriately, previous work has simply discard trajectories that do not successfully complete the task, using only successful trajectories in the fine-tuning stage. However, in tasks demanding intricate planning, reasoning, or tool usage, the volume of discarded negative samples can exceed 60%, leading to substantial data and computational resource wastage.
@@
In this paper, we explore two key questions: (1) Can LLMs learn from negative examples through fine-tuning? and (2) How can we optimize the use of negative examples to enhance agent performance? To address the first question, we fine-tune LLMs with a mix of positive and negative examples, and observe that incorporating negative examples generally yields benefits. For the second question, we introduce a negative-aware training (NAT) approach that explicitly instructs the model to differentiate between correct and incorrect interactions. Our experiments demonstrate that NAT outperforms traditional methods that solely use positive examples or naively combine positive and negative examples. In addition, we conduct extensive experiments to analyze the learned agents’ behavior when fine-tuning with negative examples. We find that LLMs implicitly learn a lot from negative samples when combined with positive ones.
Our contributions can be summarized as follows:
@@
  • We demonstrate the value of negative trajectories and introduce a negative-aware training approach, allowing LLM-based trained agents to effectively learn from both positive and negative examples.
@@
  • We validate the broad applicability and effectiveness of learning from negative examples, and show that it enables models to acquire information akin to positive examples across various tasks and prompting strategies.
Figure 1: Illustration of negative-aware training (NAT).
Figure 2: Example of positive (left) and negative (right) trajectories obtained from the GSM8K dataset.
@@ 2.1 Fine-tuning LLMs as Agents
Previous work on language agents has taken a powerful LLM as the core of the agent system without fine-tuning Sumers et al. (2023); Wu et al. (2023); Ruan et al. (2023); Zhao et al. (2023). However, LLMs are optimized to generate natural language. To make them capable of using tools and making decisions, current work typically collects trajectories generated by GPT-3.5/4, then uses only successful trajectories to fine-tune a smaller LLM. Zeng et al. (2023) collect trajectories generated by GPT-4 on AgentBench (Liu et al., 2023) tasks, and only keep samples that received the best rewards. Chen et al. (2023) collect trajectories on question answering tasks and fine-tune models with samples that correctly answered the question. Liu et al. (2024) propose a memory-enhanced agent framework and a complex filtering mechanism to collect fine-tuning datasets. Qiao et al. (2024) divide an agent into sub-agents with different functions. They then synthesize trajectories for these agents respectively. However, they still only use samples with the best rewards. Although a simple ablation study is done by Zeng et al. (2023), none of them has investigated the effectiveness of negative samples in detail. Although not directly comparable, in Table 1, we compare the methods of these papers with the ours on several main benchmarks.
2.2 Learning from Negative Trajectories
Learning from negative trajectories can be divided into prompting-only and fine-tuning-based methods. Prompting-based methods enable LLMs to summarize experiences from previous mistakes without updating parameters. Madaan et al. (2023) use LLMs to first generate an output and then refine the output iteratively, while Shinn et al. (2023) employ an evaluator to provide external feedback. Zhao et al. (2023) let the agent compare successful and unsuccessful trajectories, and extract insights based on comparison. The success of these methods relies on the quality of the evaluator used to analyze the trajectories. The performance of fine-tuning-based methods is less predictable since model weights are updated, and less work has been done on this. Li et al. (2023) propose a two-stage training paradigm to capture knowledge from negative samples. However, their method focuses on Chain-of-Thought prompts and is complex since multiple models are fine-tuned. Our work focuses on the fine-tuning-based method and is both much simpler and more effective.
@@ 3 Negative-Aware Training
3.1 Background
As shown in Figure 1, in our agent framework, the process of task resolution is delineated as follows. First, the LLM is provided with an initial prompt that outlines (a) the specific task to be addressed (for instance, solve a mathematical problem), (b) the tools that are permissible for task execution, and (c) the expected action space and output format (for example, finish[N] signifies that N is the final answer). See the example in Figure 2. Second, an instance is introduced, prompting the model to generate an initial series of “thoughts” and actions. Finally, during the interaction phase, the system executes the LLM-generated actions using the predefined tools, returns the resulting observations back to the LLM, and prompts for subsequent actions until the finished action of the task is generated, or the interaction rounds exceeds a pre-defined threshold.
Naturally, the task-solving process yields interaction trajectories between the LLM and the environment (i.e., tools in our framework). Prior research has demonstrated that fine-tuning the LLM with successful interaction trajectories can significantly enhance its problem-solving capabilities Chen et al. (2023); Zeng et al. (2023).
In this study, we propose to use not only positive examples (i.e., trajectories that lead to successful task resolution), but also negative samples (i.e., trajectories that fail to resolve the task).
@@ 3.2 Motivation
Our idea is motivated by two considerations. First, humans learn from failure, because it provides valuable lessons that contribute to personal growth and improvement. Failure is often seen as a stepping stone to success, as it offers insights into what doesn’t work and highlights areas that require change or development.
Second, even the most advanced LLMs such as GPT-4 produce unsuccessful trajectories in over 50% of attempts when acting as an agent to solve simple mathematical reasoning questions. Simply discarding these trajectories leads to a waste of computational resources, and ignores potentially valuable supervision signal.
3.3 Proposed Approach
Our method can be structured into four phases, as detailed below: (1) data collection, (2) data cleaning, (3) negative-aware reformatting, and (4) fine-tuning.
@@ 3.3 Proposed Approach
For each task, we obtain the initial questions and corresponding ground truth answers as seed data. We then use an LLM to generate trajectories three times, each with different temperatures (0.2, 0.5, and 0.7). This allows us to gather a diverse range of positive samples while experimenting with various negative sample collection strategies. In this paper, we use GPT-3.5-1106 to generate high-quality trajectories. [^2] Although GPT-4 has the potential to produce even higher quality data, we opted for GPT-3.5 due to cost considerations.
Data cleaning
We contend that certain negative samples are beneficial for fine-tuning models, while others may be harmful. Therefore, data cleaning plays a vital role in our approach. Given the variability across tasks, we adopted task-specific data-cleaning criteria and methodologies, which we detail in the experimental setup of each task.
@@
Differentiating positive samples from negative samples during the agent tuning process aids in teaching the model to discern between successful and unsuccessful outcomes. We frame this as implicit contrastive learning, to help the model learn from both success and failure. Therefore, before fine-tuning, we concatenate different prompts to positive and negative trajectories, to differentiate them to the model.
Fine-tuning
After negative-aware reformatting, we use the reformatted trajectories to fine-tune an LLM, which we call negative-aware training (NAT). The loss is computed only on the part of the text generated by the LLM, as seen in Figure 2 in the blue boxes. As an ablation, for all tasks, we also fine-tune the models on the dataset without negative-aware reformatting, where the positive and negative examples are concatenated with the same prompt. We name this training process negative-unaware training (NUT).
Inference
During inference, we prompt the fine-tuned agent using the prompt for positive examples only.
3.4 Experimental Setup
Following prior work (Chen et al., 2023; Yin et al., 2023), we conduct experiments on mathematical reasoning, multi-hop question answering, and strategic question answering tasks, using GSM8k Cobbe et al. (2021), HotpotQA Yang et al. (2018), and StrategyQA Geva et al. (2021), respectively.
Agent Framework
We use ReAct (Yao et al., 2023), which consists of interleaving “thoughts” and actions. During inference, an observation from the environment is given to the model. The model first generates a “thought”, which is plain-text reasoning based on the observation as to what it should do next. Then, it generates an action, which is what the tools use in agent scenarios. We also experiment with Chain-of-Thought (Wei et al., 2022) prompting in Section 6.
Tools
For math tasks, we design a calculator implemented by SymPy (Meurer et al., 2017), which takes a math expression as input and outputs the result. For the two question-answering tasks, we design a search tool for the Serper API. It takes a search query as input and returns the Google search results. We further re-rank the search results using MPNet (Song et al., 2020) and DPR (Karpukhin et al., 2020).
Fine-tuning Setup
We conduct experiments on LLaMA-2-Chat 7B and 13B models. All the models are fine-tuned for 2 epochs with a batch size of 64. We use a cosine scheduler with 3% of total steps as the warm-up. The maximum learning rate is set to $5\times 10^{-5}$. We train the model with 4$\times$×A100 GPUs with DeepSpeed ZeRO 3 stage (Rajbhandari et al., 2019).
@@ 3.4 Experimental Setup
4 Math Reasoning
4.1 Datasets

### v2

@@ Abstract
Large language models (LLMs) have achieved success in acting as agents, which interact with environments through tools such as search engines. However, LLMs are optimized for language generation instead of tool use during training or alignment, limiting their effectiveness as agents. To resolve this problem, previous work has first collected interaction trajectories between LLMs and environments, using only trajectories that successfully finished the task to fine-tune smaller models, making fine-tuning data scarce and acquiring it both difficult and costly. Discarding failed trajectories also leads to significant wastage of data and resources and limits the possible optimization paths during fine-tuning. In this paper, we argue that unsuccessful trajectories offer valuable insights, and LLMs can learn from these trajectories through appropriate quality control and fine-tuning strategies. By simply adding a prefix or suffix that tells the model whether to generate a successful trajectory during training, we improve model performance by a large margin on mathematical reasoning, multi-hop question answering, and strategic question answering tasks. We further analyze the inference results and find that our method provides a better trade-off between valuable information and errors in unsuccessful trajectories. To our knowledge, we are the first to demonstrate the value of negative trajectories and their application in agent-tunning scenarios. Our findings offer guidance for developing better agent-tuning methods and low-resource data usage techniques.  [^1] Code and data are available at: https://github.com/Reason-Wang/NAT.
@@ 1 Introduction
An agent is a model that has the ability to interact with environments, make decisions, and achieve predefined goals (Wooldridge, 1999). Early work used rule-based or template-based systems to complete tasks in narrow and specialized domains (Green Jr et al., 1961; Weizenbaum, 1966). Recent work has built off powerful LLMs such as GPT-4 (Achiam et al., 2023), using them as the core of an agent system to process information and make decisions (Gravitas, 2024; Yoheinakajima, 2024). This line of work has resulted in agent systems that are able to perform much more complex and general tasks.
@@
However, these agents generally rely on closed-source LLMs through paid APIs, raising concerns about cost, latency, and reproducibility. Additionally, existing LLMs were not developed for agent use cases (e.g., generating actions or calling tools), and few-shot prompting offers only limited learning support (Chen et al., 2023).
@@
Subsequent work has explored fine-tuning LLMs as agents, typically in three stages: data collection, fine-tuning, and inference (Chen et al., 2023; Zeng et al., 2023; Yin et al., 2023; Qiao et al., 2024). At the data collection stage, a powerful LLM such as GPT-4 is employed to interact with the environment, and the LLM-generated outputs and environment observations are collected as trajectories. In the fine-tuning stage, smaller models are fine-tuned using only successful trajectories. The fine-tuned models then serve as the agent’s core during inference, demonstrating enhanced tool-using and decision-making capabilities, sometimes even surpassing the performance of the original LLM (Yin et al., 2023).
@@
To ensure the agent is being optimized appropriately, previous work has simply discarded trajectories that do not successfully complete the task (i.e. negative examples), using only successful trajectories (i.e. positive examples) in the fine-tuning stage (Zeng et al., 2023; Chen et al., 2023; Qiao et al., 2024). However, in tasks demanding intricate planning, reasoning, or tool usage, the volume of discarded negative samples can exceed 60%, leading to substantial data and computational resource wastage.
@@
In this paper, we explore two key questions: (1) Can LLMs learn from these negative examples through fine-tuning? and (2) How can we optimize the use of negative examples to enhance agent performance? To address the first question, we fine-tune LLMs with a mix of positive and negative examples, and observe that incorporating negative examples generally yields benefits. For the second question, we introduce a negative-aware training (NAT) paradigm that explicitly tells the model to differentiate between correct and incorrect interactions by adding prefixes or suffixes. Our experiments demonstrate that NAT outperforms traditional methods by solely using positive examples or naively combining positive and negative ones, enabling better fine-tuning for low-resource data. In addition, we conduct extensive experiments to analyze the learned agents’ behavior after fine-tuning with negative examples. Our contributions can be summarized as follows:
@@
  • We demonstrate the value of negative trajectories and introduce a negative-aware training paradigm, allowing LLM-based trained agents to effectively learn from both positive and negative examples. To the best of our knowledge, we are the first to utilize negative examples in agent training.
@@
  • We validate the broad applicability and effectiveness of learning from negative examples, and show that NAT enables models to acquire information akin to positive examples across various tasks and prompting strategies.
  • We find NAT works by providing a better trade-off between useful information and errors in negative examples.
Figure 1: An overview of previous methods and our NAT paradigm. (a) Data collection, where interactions between LLMs and environments (tools) are collected. (b) Data processing, where previous methods simply filter out negative examples, while we reformat trajectories by adding prompts to task queries based on whether they are positive or negative. (c) An example of reformated positive and negative trajectories. We omit the system prompts here.
@@ 2.1 Fine-tuning LLMs as Agents
Previous work on language agents has taken a powerful LLM as the core of the agent system without fine-tuning Sumers et al. (2023); Wu et al. (2023); Ruan et al. (2023); Zhao et al. (2023). However, LLMs are optimized to generate natural language. To make them capable of using tools and making decisions, current work typically collects trajectories generated by GPT-3.5/4, then uses these trajectories to fine-tune a smaller LLM (Zeng et al., 2023; Chen et al., 2023; Qiao et al., 2024; Chen et al., 2024; Zhang et al., 2024; Zhou et al., 2024). Zeng et al. (2023) collect trajectories generated by GPT-4 on AgentBench (Liu et al., 2023b) tasks, and only keep samples that receive the best rewards. Chen et al. (2023) collect trajectories on question answering tasks and fine-tune models with samples that correctly answer the question. Liu et al. (2024) propose a memory-enhanced agent framework and a complex filtering mechanism to collect fine-tuning datasets. Qiao et al. (2024) divide an agent into sub-agents with different functions. They then synthesize trajectories for the respective agents. However, they still only use samples with the best rewards. A simple ablation study was done by Zeng et al. (2023). However, none of this work has investigated the effectiveness of negative samples in detail. Although not directly comparable, in Table 1, we provide the results of these methods and ours on several benchmarks for reference.
2.2 Learning from Negative Results
Learning from negative results can be divided into prompt-based and fine-tuning-based methods. Prompt-based methods enable LLMs to summarize experiences from previous mistakes without updating parameters. Madaan et al. (2023) use LLMs to first generate an output and then refine the output iteratively, while Shinn et al. (2023) employ an evaluator to provide external feedback. Zhao et al. (2023) let the agent compare successful and unsuccessful trajectories, and extract insights based on comparison. The success of these methods relies on the quality of the evaluator used to analyze the trajectories. The performance of fine-tuning-based methods is less predictable since model weights are updated, and less work has been done on this. Li et al. (2023) propose a two-stage training paradigm to capture knowledge from negative samples. However, their method focuses on Chain-of-Thought prompts and is complex since multiple models are fine-tuned. Our work focuses on fine-tuning LLMs as agents and is much simpler and more effective.
@@ 3 Negative-Aware Training
We first illustrate our motivation for using NAT and then describe our agent framework. We then introduce the whole pipeline of NAT, including data collection, negative-aware reformatting (which is the core part of our method that differentiates it from others), fine-tuning, and inference. Figure 1 outlines previous methods and our NAT paradigm.
@@ 3.1 Motivation
Our idea is motivated by two considerations. First, humans learn from mistakes and failures. Failure is often seen as a stepping stone to success, as it offers insights into what does not work and highlights areas that require change or development. We believe that powerful LLMs can also learn these valuable lessons from unsuccessful trajectories.
Second, it is generally hard for humans to compare and learn from experiences when they do not know which experience is successful. Fine-tuning approaches generally treat all examples equally, and LLMs may learn unwanted errors from negative trajectories if negative examples are incorporated directly. Therefore, we add a prefix or suffix to the query to differentiate positive and negative examples, explicitly telling the model whether the following trajectories they learn are correct.
3.2 Agent Framework
Prompting Strategy
As shown in Figure 1, in our agent framework, the process of task resolution is delineated as follows. First, the LLM is provided with a system prompt that outlines (a) the specific task to be addressed (for instance, ‘‘solve a mathematical problem’’), (b) the tools that are permissible for task execution, and (c) the expected action space and output format (for example, finish[N] signifies that N is the final answer). We do not provide system prompts in Figure 1, for simplicity. Second, a query instance is introduced. We prompt the model to answer the query in the ReAct (Yao et al., 2023) format, which consists of reasoning texts (referred to as “thoughts”) and “actions”. Finally, during the interaction phase, the system executes the LLM-generated actions using the predefined tools, returns the resulting observations back to the LLM, and prompts for subsequent actions until the finish action of the task is generated, or the interaction rounds exceed a pre-defined threshold. Naturally, the task-solving process yields interaction trajectories between the LLM and the environment (i.e., tools in our framework).
Tools
For math tasks, we design a calculator implemented by SymPy (Meurer et al., 2017), which takes a math expression as input and outputs the result. For the two question-answering tasks, we design a search tool with the Serper  [^2] https://serper.dev/ API. It takes a search query as input and returns the Google search results. We further re-rank the search results using MPNet (Song et al., 2020) and DPR (Karpukhin et al., 2020).
3.3 The Whole Pipeline
We introduce the whole pipeline of our negative-aware training paradigm here, where negative-aware reformatting is the core part of the paradigm that enables better agent tuning.
@@ 3.3 The Whole Pipeline
For each task, we obtain the initial questions and corresponding ground truth answers as seed data. We then use GPT-3.5333We use GPT-3.5-1106 version. Although GPT-4 has the potential to produce even higher quality data, we opted for GPT-3.5 due to cost considerations. to generate trajectories three times, each with different temperatures (0.2, 0.5, and 0.7). This allows us to gather a diverse range of positive and negative samples. By comparing predicted answers and ground truth answers, we can label each trajectory as positive or negative.
@@
Differentiating positive samples from negative samples during the agent tuning process aids in teaching the model to discern between successful and unsuccessful outcomes. We append a string suffix to tell the model whether the training sample is positive or negative. For positive samples, we append ‘‘Please generate a solution that **correctly** answers the question.’’ For negative samples, we append ‘‘Please generate a solution that **incorrectly** answers the question.’’ Unless explicitly stated, we use this in experiments. We also experimented with other reformatting strategies. [^4] The actual prompts that we use in our experiments are slightly more complex than those provided here.
Fine-tuning and Inference
We use the reformatted trajectories to fine-tune LLMs. The loss is computed only on the part of the text generated by the LLM, which is similar to fine-tuning a chat model (Zheng et al., 2023). During inference, we prompt the fine-tuned agent using the prompt for positive examples only.
@@
Strategy 7B 13B
EM F1 EM F1
Vanilla 27.44 36.41 27.04 36.95
NUT 28.04 40.96 28.24 42.47
NAT 28.80 41.37 28.44 42.45
NAT-2 29.76 42.51 29.60 43.29
Table 3: Results of LLaMA-2-7B and 13B on HotpotQA. All results are reported as the mean score of 5 runs. For the 7B model, we report results using 1,500 negative samples; for the 13B models, we use 2k negative samples. NAT-2 means we divide negative samples into 2 classes based on quality. We discuss this in Section 6.1.
Strategy 7B 13B
Vanilla 55.40 53.40
NUT 62.40 61.80
NAT 65.80 64.60
4 Experiments
4.1 Experimental Setup
Datasets
We conduct experiments on mathematical reasoning, multi-hop question answering, and strategic question answering tasks. For math, we collect trajectories using GSM8k Cobbe et al. (2021) as seed data and test the performance on GSM8k, ASDiv (Miao et al., 2020), SVAMP (Patel et al., 2021), and MultiArith (Roy and Roth, 2015). For question answering, we collect trajectories and test the performance on HotpotQA Yang et al. (2018) and StrategyQA Geva et al. (2021), respectively. More details are provided in Appendix B.
Baselines
We primarily compare NAT with two baselines. The Vanilla setting uses positive examples only to fine-tune LLMs. This is what previous work (Zeng et al., 2023; Chen et al., 2023; Qiao et al., 2024; Liu et al., 2024) has done. The second setting includes negative examples without adding any prefix or suffix, which we call Negative-Unaware Training (NUT).
Fine-tuning Setup
We conduct experiments on LLaMA-2-Chat 7B and 13B models (Touvron et al., 2023). All the models are fine-tuned for 2 epochs with a batch size of 64. We use a cosine scheduler with 3% of total steps as the warm-up. The maximum learning rate is set to $5\times 10^{-5}$. We train the model with 4$\times$×A100 GPUs with DeepSpeed ZeRO 3 stage (Rajbhandari et al., 2019).
4.2 Results
Math
Table 2 presents the overall results of the math tasks, from which we observe: (1) Incorporating negative examples can improve model performance. (2) Models with negative-aware training (NAT) not only outperform the corresponding model trained only on positive examples (Vanilla), but also beat the same model trained by directly incorporating negative examples (NUT); and (3) The improvement of NAT is more substantial when there are fewer positive examples or the model is smaller. Specifically, NAT achieves an $8.74\%$ improvement when using a 7B model with 2k positive examples, and a $0.52\%$ improvement when using a 13B model with 5k positive examples. This highlights the value of NAT in data-scarce scenarios, which is common for agent tuning.
It is worth noting that previous work (Zeng et al., 2023) has shown that including negative examples harms model performance. We believe this does not contradict our findings: as we discuss in Sections 5.2 and 5.1, performance is determined by both the quantity and quality of the negative data.
Question Answering
Tables 3 and 4 show the results on HotpotQA and StrategyQA. Here, NAT-2 is a variant of NAT, where we divide negative data into two classes and use different prompts for each, as detailed in Section 6.1. On HotpotQA, NAT-2 improves performance by more than $2\%$ in EM and $6\%$ in f1 score compared to no negative samples. Compared with NUT, NAT is still about $1\%$ better on EM and f1. On StrategyQA, NAT achieves more than $8\%$ and about $3\%$ improvements compared to no negative samples and NUT, respectively. This suggests that our method is also effective for question-answering tasks.
5 Analysis
Tables 2, 3 and 4 showcase the capability of LLMs to learn from negative examples when fine-tuned to function as agents. Here, we delve into various factors that could influence the effectiveness of negative-aware training. Specifically, we seek to address the following questions: (1) Given a fixed number of positive examples, how much negative data should be used? (2) What insights does the model gain from negative trajectories? (3) Are all types of negative examples beneficial? and (4) What factors contribute to negative-aware training (NAT) outperforming negative-unaware training (NUT)? Since only the math task contains enough data for our experiments, the analysis is done on the math task.
5.1 Impact of Training Sample Quantity
Our initial analysis focuses on the influence of negative sample quantity. We maintain a constant number of positive samples at 2k and 5k, while adjusting the negative samples from 0 to 12k. The results, depicted in Figure 2, illustrate the relationship between the quantity of negative data and average math task performance. We observe a performance enhancement with an increase in negative data, which plateaus when the volume of negative samples is about 11k in both cases. Due to data availability, we did not experiment with more negatives.
Based on insights from Table 2 and Figure 2, we hypothesize that the ideal ratio of negative samples is not fixed. Instead, it is influenced by two main factors: (1) the number of positive samples, as the improvements are larger for fewer positive samples; and (2) the intrinsic quality of the negative samples.
Regarding the first point, we hypothesize that the marginal utility of negative samples diminishes as the quantity of positive samples increases. To test this point, we maintain a constant number of negative samples while varying the quantity of positive samples from 0 to 5k. As depicted in Figure 3, there is a diminishing return on the performance added by negative samples as the count of positive samples rises. For the second point, we investigate the effects of negative data quality in Section 5.2.
Data GSM8K ASDiv SVAMP MArith Avg.
2K positive samples
Vanilla 35.63 60.55 47.40 80.03 55.90
NAT-low 32.98 58.72 47.60 71.64 52.74
NAT-high 46.93 66.93 60.80 83.89 64.64
5K positive samples
Vanilla 45.87 68.12 58.80 83.89 64.17
NAT-low 38.59 62.28 52.50 78.52 57.97
NAT-high 49.05 68.66 64.40 87.58 67.42
5.2 Data Quality Matters
We sourced negative data from various models to investigate the impact of negative data quality in NAT. Specifically, we consider the data from GPT-3.5 as high-quality examples. In contrast, we generated 10k negative examples using a fine-tuned LLaMA-2-7B (Touvron et al., 2023) model to represent low-quality data. For experiments, we paired 2k positive examples with 10k negative examples. The outcomes presented in Table 5 underscore the critical role of data quality in NAT. In the 2k positive sample setting, the improvement is $-3.16$ for low quality compared to $+8.74$ for high-quality negative examples. Similarly in the 5k positive sample setting, the improvements are $-6.20$ and $+3.25$, respectively.
5.3 What does the Model Learn with NAT?
Learning Reasoning Rather Than Acting
The learnable part in trajectories are thoughts and actions, where thoughts are the reasoning on the current situation and planning for what to do next. Actions are which tool to call and the input to the tool. We analyze the trajectories of the GSM8K Cobbe et al. (2021) test set generated by LLaMA-2-7B trained with positive examples (Vanilla), NUT, and NAT respectively. Table 6 shows the accuracy, action error (the percent of incorrectly calling a tool), and average turns (the average number of steps needed to solve a question). Incorporating negative examples also introduces more action errors, which can result in fine-tuned models with more action errors compared to Vanilla. However, after incorporating negative examples, both the accuracy of NUT and NAT increase. This indicates that negative examples mainly work by teaching models with better “thoughts” (i.e. reasoning and planning). Compared to NUT, NAT achieves significantly fewer action errors and, therefore, better accuracy. This demonstrates that our method works by providing a better trade-off between better “thoughts” and more action errors.
Strategy Accuracy Action Error #Avg. Turns
Vanilla 35.63 3.58% 3.12
NUT 44.43 10.47% 3.92
NAT 46.93 7.90% 3.71
Table 6: Accuracy, action error rate, and number of average turns for models with different training strategies. For Vanilla, the action error in training data is 4.01%. For NUT and NAT, it is 15.33%.
Negative Samples Play a Similar Role as Positive Samples
To further explore whether models learn from negative trajectories in the same manner as they learn from positive trajectories, we randomly sample 100 successful trajectories from the training set (as a dev set) and measure the perplexity of models trained with 2k positive examples (not overlapping with the dev set) and varying numbers of negative examples. Figure 4 shows the change in perplexity as the number of negative data increases. The perplexity decreases as more negative data is included, which indicates the model learns to fit successful trajectories with knowledge from failed trajectories. However, this curve seems to be horizontal at the end, and there is still a large gap between the curve with 4k and 5k positives, which shows that some properties or knowledge from successful trajectories can never be learned from failed trajectories.
Positive Negative Average
Correct Incorrect 63.55
Incorrect Correct 63.33
Good Bad 63.91
A B 63.15
Random string 1 Random string 2 64.04
Table 7: Results for models trained on prompts with and without interpretability. Strings in the Positive/Negative column represent prompts (prefixes or suffixes) we use for positive/negative trajectories.
5.4 Selection of Added Prompts
It has variously been shown that prompts are vital for LLM performance (Brown et al., 2020; Liu et al., 2023a; Sclar et al., 2024). Here, we explore the interpretability of added prompts. More specifically, does the content of the prompt enable LLMs to learn differently from successful and failed trajectories, or simply differentiate these trajectories? We propose two sets of prompts. One set is prompts with interpretability, such as having the model generate a correct or incorrect trajectory. Another set is prompts without interpretability. For example, different letters can be added as prefixes for queries. Table 7 shows the results of models trained with interpretable and uninterpretable prompts. Different prompts do not show a large difference in performance, indicating that the performance boost of NAT comes from simply differentiating positive and negative data.
Strategy GSM8K ASDiv SVAMP MArith Avg
Vanilla 29.04 55.26 45.60 80.87 52.69
NUT 33.50 61.69 52.20 86.41 58.45
NAT 36.24 61.10 53.90 86.24 59.37
6 Applications
In this section, we explore some other applications with our NAT method, showcasing its potential and broad applicability.
6.1 Fine-grained NAT
Different negative trajectories contain different degrees of errors. Intuitively, this information also helps models to learn. Therefore, we further propose fine-grained NAT, which divides negative trajectories into different groups based on their quality. During training, different groups will be reformatted with different prompts. For HotpotQA, in addition to the EM score, each trajectory has an f1 score, measuring the overlap between the predicted and gold answers. We take this as a fine-grained measurement of data quality, where a trajectory with a higher f1 score has better quality. In this way, we can differentiate trajectories based on quality by assigning different prompts. For example, the trajectory is prepended ‘‘almost wrong’’ if its f1 score is smaller than 0.1, and another trajectory is ‘‘mostly correct’’ with an f1 score of 0.9. We denote this NAT with different prompting strategies as NAT-$k$, where $k$ represents how many classes we divide the negative data into based on quality.
Fine-grained NAT learns more from negative samples
For NAT-2, we take trajectories with f1 scores equal to 1.0 as positive and assign different prompts for trajectories with f1 scores less than 0.4 and with f1 scores greater than 0.4 less than 1.0. It can be seen from Table 3 that the NAT-2 consistently outperforms NAT in all settings, indicating that negative samples associated with finer-grained info are more informative. We investigate how the performance changes with differing numbers of negative samples in Figure 5. When adding negative samples, the performance increases by a margin, consistent with Table 3. However, NAT achieves the best performance with only 500 negative samples, and its performance decreases when adding more negative samples. NAT-2, on the other hand, achieves the best performance with 1k negative samples, consistent with our hypothesis that fine-grained NAT is more beneficial to model training.
6.2 Chain-of-Thought Prompting
So far we have conducted all experiments on agent scenarios with the ReAct (Yao et al., 2023) prompting strategy. In this section, we conduct preliminary experiments to explore whether NAT works well with Chain-of-Thought (CoT) prompting Wei et al. (2022). The key difference is that the agent takes an action and receives an observation from the environment iteratively, while CoT generates reasoning steps without taking actions or receiving observations.
We use GPT-3.5-0125 to generate CoT reasoning steps with three in-context learning (Brown et al., 2020) examples on the GSM8k dataset. We then train the model with NAT. Table 8 shows the results with CoT prompting. NAT achieves a $6.68\%$ improvement compared to no negative data training (Vanilla). NAT is still about $1\%$ higher compared to directly including negative samples (NUT). The results demonstrate that NAT is also applicable and effective for CoT training, showing its broad applicability.
7 Conclusion
In this paper, we first demonstrated that LLMs can learn from failures when fine-tuned as an agent. On the basis of this finding, we propose NAT, a simple and effective method for integrating failed trajectories in fine-tuning agents. We conduct experiments on math and question-answering tasks, and show the superior performance of our method compared to training directly with positive or negative trajectories across tasks and model sizes.
We are the first to demonstrate the value of negative trajectories in fine-tuning LLMs as agents. Our analysis finds that the quality of negative data is the key to the success of NAT, and models learn similar knowledge to that of positive data (which is much more expensive to attain). NAT is superior because it better utilizes valuable information while restraining learning errors from negative examples.
Negative-aware training is designed to be both agent-agnostic and reasoning strategy-agnostic, making it compatible with various agent strategies, including self-refinement (Madaan et al., 2023), reflection (Shinn et al., 2023), and other agent-tuning frameworks. At the end of this paper, we demonstrated the effectiveness of NAT on Chain-of-Thought (COT) reasoning in mathematical tasks. Moving forward, we aim to assess the applicability and effectiveness of NAT across a broader spectrum of agent frameworks, strategies, and tasks.
8 Limitations
Although we have conducted extensive experiments to illustrate the effectiveness of our method, there are still limitations. First, similar to previous work in agent tuning, our approach requires the ground truth label of the data, which limits its application. Second, we do not experiment with fine-tuning our method on more diverse and powerful models (e.g. GPT-3.5) due to time and budget limits.
Appendix A Example
Figure 6 shows examples trajectories generated by GPT-3.5. The first turn of each trajectory is the system prompt. Figure 7 shows example inference results of models trained with different settings for the same query.
Figure 6: Example trajectories of GSM8k, HotpotQA, and StrategyQA generated by GPT-3.5.
Figure 7: Inference results of models trained with different settings, where Vanilla model gets wrong because of reasoning error, NUT model gets wrong because of action error, and NAT gets correct.
Appendix B Datasets